OpenStudy (anonymous):

A balloon has a circumference of 28 in. Use the circumference to approximate the surface area of the balloon to the nearest square inch.

OpenStudy (abmon98):

Circumference=D(Diameter)*Pi D=28/Pi=8.9126 In order to find the balloons radius we divide our diameter by 2 so we get 4.4563 we use the formula 4*pi*r^2 4*pi*(4.4563)^2